What is the acceleration due to gravity on a planet with a mass of 4.25x10^23 kg and a mean equatorial radius of 5.54x10^6 m?

(G=6.67x10^-11 N.m^2 / Kg^2)
1
Expert's answer
2020-10-13T09:30:09-0400

"g=F/m=G\frac{Mm}{R^2}/m=G\frac{M}{R^2}="


"=6.67\cdot10^{-11}\cdot \frac{4.25\cdot10^{23}}{(5.54\cdot10^6)^2}=0.92(m/s^2)"